# HandProxy: Expanding the Affordances of Speech Interfaces in Immersive Environments with a Virtual Proxy Hand

### v1

@@ 4.1. Speech Understanding
The speech understanding module is responsible for recognizing and interpreting the user’s command. Upon receiving a spoken command, the system first determines whether it is an interaction command or a random speech input, checks if it is ambiguous or interpretable, and then decomposes it into structured, system-interpretable instructions for the hand control module to execute.
For speech recognition, HandProxy uses the Google Speech-to-Text API (Google, 2024), which is configured to return intermediate results for real-time streaming recognition. The speech input is segmented into individual sentences, determined by terminal punctuation on the final results returned from the API. Each sentence is then treated as a command and is added to the command queue to be processed in later stages.
For command interpretation, HandProxy uses a multi-level approach for command processing. Similar to the reactive and deliberative layers in robotics (Murphy, 2019), the system uses keyword matching and directly responds to high-priority commands, with LLM-based processing for more complex inputs. Specifically, as soon as intermediate results are received from the speech recognition service, HandProxy checks predefined keywords and synonyms related to temporal controls as discussed in section 3. These commands — such as “stop,” “hold,” or “undo” — are time-sensitive and need to be executed with minimal delay. If a matching keyword is detected, the system triggers the corresponding temporal control immediately. Otherwise, the system retrieves the earliest available command from the command queue and processes it using GPT-4o (OpenAI, 2024). The model is guided by a structured system prompt and is provided with a list of supported control categories from section 3, which is listed below:
@@
This structured approach ensures that the LLM interprets user commands accurately and maps them to the appropriate hand interactions, which includes:
@@
Using this information, the LLM is instructed to (i) correct potential speech recognition errors (e.g., power bottom $\rightarrow$ power button), which has been used to improve speech recognition results (Yang et al., 2023), (ii) decompose high-level commands, if any, into executable steps (e.g., put the cube into the basket $\rightarrow$ grab the cube, move to the basket, and release), and (iii) parse individual steps as a list of instructions as json. Also, since people tend to use a surprisingly great variety of words to refer to the same thing (Furnas et al., 1987), the system prompt instructs the model to generalize beyond exact words provided in the supported instructions to accommodate diverse ways users may phrase their commands. For example, commands with similar intent are mapped to the same control action (e.g., “pinch” = “tap index and thumb,” “pink circular fruit” = “peach”). If a given input is not relevant — meaning it does not appear to be an interaction command — the model is instructed to ignore it and return an empty list instead of attempting to interpret unrelated speech. Additionally, user commands are saved within a command list, allowing the system to refer back to previous inputs when necessary (e.g., resolving references to earlier commands).
With this approach, a user’s command can be converted into system-executable control instructions. For example, given the command “pinch the cube,” the system generates the following instruction:
@@
By dynamically generating one or more executable steps, HandProxy provides users with the flexibility to issue both low-level direct commands and high-level intent-based commands, which provides an adaptive and natural interaction experience.
The system latency is composed of two main components: speech recognition and processing latency, and command processing latency. Speech recognition and processing latency measures the time from just before an audio chunk is sent via an API request to when the recognized text is split into commands. Based on 100 requests, the average latency for this process is 0.18 seconds, with a standard deviation of 0.07 seconds. The command processing time is the time it takes for the system to interpret the command and generate a sequence of corresponding hand pose keypoints. We measured such latency for both commands that matched high-priority keywords, and those processed by LLMs. Through a test of 100 commands that match the high priority keywords, all commands were executed within 0.001 seconds. Through a test of 100 requests that include both simple (e.g., requires one step), complex (e.g., needs to be decomposed into multiple steps), and irrelevant commands (e.g., not a valid interaction instruction), the average response time was 1.39 seconds, with an standard deviation of 0.81 seconds.
[…]
@@ 4.2. Hand Control
The hand control module executes a given list of control instructions to generate a sequence of hand pose data to control the virtual hand. It maintains a hand state manager, which tracks key hand properties such as the current hand pose, wrist position, playback status (e.g., whether an action is in progress), and the sequence of upcoming hand poses. When a new instruction arrives, the hand control module activates the relevant control sub-modules discussed in section 3 to update the hand pose data. Next, we describe each control module in detail.
@@ 4.2.1. Gesture Control
The gesture control module generates a sequence of hand pose data to represent the required gesture. These poses are retrieved from the FPHAB (Garcia-Hernando et al., 2018) dataset or captured using LeapMotion, but is extensible to gesture datasets that include the 2.5D coordinates (x_ratio, y_ratio, z_depth) of 21 hand joints. The gestures supported in the current system includes grab, pinch, point, push, swipe, cut, punch, squeeze, thumb up/down, and open hand (release).
To enrich data with additional gesture information, each gesture is manually annotated with additional details about the gesture’s object dependency and different gesture stages (McNeill, 1992). While it is not a comprehensive list of gesture metadata, it provides additional information that could benefit the gesture control. Some key attributes include:
@@
    Segments: Defines the start and end frames for different gesture phases, including preparation, stroke, and retraction. This segmentation supports additional features such as gesture holding.
Combining all these data together, when given a gesture command, the module returns either a single hand pose (for static gestures) or a sequence of hand poses (for dynamic gestures). These pose data sequences are then forwarded for processing in subsequent stages.
@@ 4.2.2. Target Control
The target control module handles object-dependent hand interactions by identifying and aligning the virtual hand with the intended target. It first resolves the target object referenced using positional constraints, such as “below,” “above,” “to the left of,” “to the right of,” “in front of,” “behind,” “closest,” “farthest,” “first,” “last,” “on the left,” “in the middle,” and “on the right.” These constraints are resolved by retrieving the current positions of all virtual objects and calculating their relative spatial relationships.
Once the target object has been determined, the module updates its position by querying the virtual environment for the latest object data. It then calculates a path to move the hand from its current position to the target object. This calculation also accounts for the interacting joint (as discussed in Gesture Control) to ensure precise alignment. For example, if the command is to point at a button, the system uses the interacting joint defined in the gesture metadata to set the index fingertip to align with the center of the button. If the task is to push a cube, the system adjusts the middle finger metacarpophalangeal (MCP) joint to match the cube’s position. By dynamically modifying the original hand pose sequence, the target control module ensures that gestures interact with objects in the correct fashion. The updated hand pose data is then passed to the spatial control module for additional modifications.
@@ 4.2.3. Spatial Control
The spatial control module manages the movement and rotation of the virtual hand. It supports translational movements in six directions: “up,” “down,” “left,” “right,” “forward,” and “backward.” Additionally, it enables rotational adjustments, including “pan left,” “pan right,” “roll left,” “roll right,” “tilt up,” and “tilt down.” Specifically, translational movements are applied by adding a speed vector to the hand pose data, while rotational adjustments are calculated by rotating the joint coordinates along the estimated hand axis using the vector of middle MCP to index MCP, middle MCP to wrist, and the vector perpendicular to these two vectors.
@@ 4.2.4. Temporal Control
Temporal control specifies the timing and state changes of the hand interaction, which includes “stop”, “continue”, “faster”, “slower”, “undo_step”, “redo_step”, and “hold.” Hand pose data generated from the previous stages is sent to the hand state manager, which keeps a playback speed (frames per second, fps) and playback state (stop, play) that can be adjusted based on temporal control commands. For example, the stop and continue instructions directly change the playback state, and faster and slower change the playback speed accordingly. For the hold command, specifically when it is used to hold a specific interacting gesture (e.g., as in grab and hold the cube), the temporal control adds a key frame at the interacting frame (as defined in gesture control) and pauses the playback at that frame to hold the gesture.
Additionally, the hand state manager maintains a history of final hand poses after each executed command. This enables “undo” and “redo” functionality, allowing users to revert to previous hand states or redo an action if needed. If an error occurs, the system can quickly restore the last hand state, providing a flexible interaction experience.
@@ 4.3. Visualization and Feedback
To enhance system transparency and clarify potential ambiguities in user commands, HandProxy incorporates various visual feedback mechanisms that are overlaid on the user’s viewport. Figure 5 shows examples of each feedback type. To ensure users understand how their commands are processed, the system displays both the recognized (full input) and active (which part of the input is being executed) commands. If a command is irrelevant (e.g., it does not correspond to a valid hand interaction) or cannot be interpreted, the system highlights an error message in the recognized input and prompts the user to rephrase the command. For cases where the user refers to an ambiguous object (i.e., multiple similar objects exist in the environment), the system provides disambiguation hints by overlaying numbered labels on each possible target. The user can then restate the command while specifying the correct object by its assigned number. To further increase transparency in hand movement execution, the system visualizes the movement path. As shown in Figure 5d, a sequence of arrows appears along the trajectory, indicating the direction in which the hand is moving. If the command involves multiple steps, the full movement sequence is previewed to give users an overview of how the hand will execute the task. These visualizations provide users with greater visibility of the system to bridge the gulf of evaluation, improving both command accuracy and interaction efficiency while reducing errors and misinterpretations.
@@ 4.4. Design Iterations
The HandProxy system was iteratively designed and developed through pilot studies with 13 participants. Here, we summarize key aspects from these iterations, with the goal of providing additional rationale behind our design choices and insights into user preferences.
@@ 4.4.1. Enhancing the Speech Interface
The speech interface is a core component of HandProxy. In our initial design, we followed a traditional turn-based interaction approach, similar to off-the-shelf speech interfaces (app, 2024c), where users issue a command, wait for execution, and then issue the next command. However, our pilot studies revealed that this approach was inefficient for hand control. Many participants preferred to issue multiple commands continuously, describing an entire sequence of actions in one go. The forced pauses between commands disrupted the interaction flow, making it less natural. Additionally, these pauses made it difficult for users to issue commands to override current executions, such as stop or undo, as users had to wait for the previous command to finish before issuing a correction.
As a result, we redesigned the speech interface to support continuous input, allowing simultaneous speech recognition and action execution. This change also motivated the parallelized system architecture, where each module runs in a separate thread, ensuring that speech processing, command interpretation, and execution run without blocking other processes.
Additionally, we observed variations in user preferences regarding the level of detail in commands. Some participants preferred step-by-step instructions, such as “move left,” “move up,” “grab,” while others issued higher-level commands, such as “put the cube into the basket,” expecting the system to infer intermediate steps. This variability highlighted the limitations of rule-based systems, which rely heavily on predefined commands and synonyms. To accommodate flexible and diverse speech patterns, we integrated LLMs (GPT-4o) into the speech interface, leveraging their reasoning capabilities to interpret user intent dynamically.
@@ 4.4.2. Integrating Contextual Information
Throughout our iterations, we observed that participants naturally used descriptive references to identify target objects, including shape, appearance, and relative location. For example, participants used commands such as the “green ball” or “the object next to the button.” This demonstrated the need for contextual understanding, which motivated our decision to incorporate object metadata into the system.
@@ 4.4.3. Improving System Transparency
Speech recognition errors are a common challenge, and pilot study users expressed the need for visibility into how their commands were interpreted. This feedback motivated several transparency-focused design choices, including showing the recognized results and visualizing the path of the expected hand movement. These features ensure that users can intervene in real-time, using commands like stop or redo when necessary, ultimately improving interaction reliability and usability.
@@ 5. User Study
We conducted a user study with 20 participants to evaluate the HandProxy system and gather insights on usage patterns and feedback. Specifically, we aim to answer the following research questions:
@@ 5.1. Participants
We recruited 20 participants (9 female, 11 male, age 18-31) from the student population at our institution. According to a self-reported survey, 3 participants had no prior experience in immersive environments (e.g., VR/AR), 13 were beginners, and 4 were intermediate users. For experience with hand-gesture-controlled interfaces, 5 participants reported no experience, 12 were beginners, and 3 were intermediate users, which included devices with hand-tracking functionality, such as Apple Vision Pro, Meta Quest, Oculus Rift, and Xbox Kinect. Regarding speech interfaces, 1 participant had no experience, 5 were beginners, 12 were intermediate users, and 2 were experts. Examples of speech systems included virtual assistants like Siri or Alexa, ChatGPT voice mode, and speech input on TV remotes. The study was approved by the IRB, and participants received \$30 gift cards as compensation.
@@ 5.2. Apparatus and Procedure
The study sessions were held in a small meeting room. Participants sat in front of a TV connected to a workstation laptop with i9-12900H CPU, which displayed the virtual environment. This setup is to provide an equivalent experience for all participants. As shown in Figure 6, the virtual environment is designed to have different 3D objects and UI widgets, supporting a variety of interaction tasks. Each study session took 2 hours.
@@
    Hand interactions: participants were asked to perform a wide range of interaction tasks for virtual environment (Monteiro et al., 2021), including object selection, manipulation, transformation. The tasks included object dependent gestures (grab {apple, peach, blue cube}, press {confirm, minimize, power} button, pinch {blue cube, volume slider, resize button}, push {confirm, like, power}) and disambiguation (grab the {left, middle, right} watermelon among the 3 watermelons). A Likert scale questionnaire on system usability was given at the end.
@@
    Movement and rotation control: this included hand movement (move {left, up, forward}), object movement (put the {apple, peach, blue cube} into the basket, rotation (turn the brightness knob {clockwise, counterclockwise}). Participants were also asked to perform tasks that required timing control. This included grab the {apple, peach, cube} and move to the left according to a light signal (green = move, red = stop), and to press and hold the {power, like, confirm} button. A Likert scale questionnaire on system usability was given at the end.
@@
5.3. Data Collection and Analysis

### v2

@@ 4.2. Speech Understanding
The speech understanding module recognizes and interprets the user’s command. It determines whether the speech command is an interaction command or a random speech, checks if it is ambiguous or interpretable, and decomposes it into structured, system-interpretable instructions for the hand control module to execute.
For speech recognition, HandProxy uses the Google Speech-to-Text API, which is configured to return intermediate results for real-time streaming recognition. The speech input is segmented into individual sentences, determined by terminal punctuation on the final results returned from the API. Each sentence is then treated as a command and is added to the command queue for interpretation.
For command interpretation, HandProxy uses a multi-level approach for command processing. Similar to the reactive and deliberative layers in robotics (Murphy, 2019), the system uses keyword matching and directly responds to high-priority commands, with LLM-based processing for more complex inputs. Specifically, as soon as intermediate results are received from the speech recognition service, HandProxy checks predefined keywords and synonyms related to temporal controls as discussed in section 3. These commands — such as “stop,” “hold,” or “undo” — are time-sensitive and need to be executed with minimal delay. If a matching is detected, the system triggers the corresponding temporal control immediately. Otherwise, the system retrieves the earliest available command from the command queue and processes it using GPT-4o. The model is guided by a structured system prompt and is provided with a list of supported control categories from section 3, which is listed below:
@@
The LLM is used to (i) correct potential speech recognition errors (e.g., power bottom $\rightarrow$ power button), as used in other works like (Yang et al., 2023), (ii) decompose high-level commands, if any, into executable steps (e.g., put the cube into the basket $\rightarrow$ grab the cube, move to the basket, and release), and (iii) parse individual steps as a list of instructions as json. Also, since people tend to use a surprisingly great variety of words to refer to the same thing (Furnas et al., 1987), the system prompt instructs the model to generalize beyond exact words provided in the supported instructions to accommodate diverse ways users may phrase their commands. For example, commands with similar intent are mapped to the same control action (e.g., “pinch” = “tap index and thumb,” “pink circular fruit” = “peach”). If a given input is not relevant — meaning it does not appear to be an interaction command — the model is instructed to ignore it and return an empty list instead of attempting to interpret unrelated speech. Additionally, user commands are saved within a command list, allowing the system to refer back to previous inputs when necessary (e.g., resolving references to earlier commands).
This approach ensures that the LLM interprets user commands and maps them to the appropriate hand interactions, including:
@@
At the end, the speech understanding module outputs a json object following this format. For example, given the command “pinch the cube,” the system generates the following instruction:
@@
By dynamically generating one or more executable steps, HandProxy lets users issue both low-level direct commands and high-level intent-based commands, which provides an adaptive and natural interaction experience.
@@ 4.3. Hand Control
The hand control module executes instructions from the json to generate a sequence of hand pose data to control the virtual hand. It maintains a hand state manager, which tracks key hand properties such as the current hand pose, wrist position, playback status (e.g., whether an action is in progress), and the sequence of upcoming hand poses. When a new instruction arrives, the hand control module activates the relevant control sub-modules discussed in section 3 to update the hand pose data. Next, we describe each control module in detail.
@@ 4.3.1. Gesture Control
The gesture control module generates a sequence of hand pose data to represent the required gesture, which are retrieved from the FPHAB (Garcia-Hernando et al., 2018) dataset or captured using LeapMotion, but is extensible to gesture datasets that include the 2.5D coordinates (x_ratio, y_ratio, z_depth) of 21 hand joints. Currently supported gestures include grab, pinch, point, push, swipe, cut, punch, squeeze, thumb up/down, and open hand (release).
To enrich data with additional gesture information, each gesture is manually annotated with details about the gesture’s object dependency and different gesture stages (McNeill, 1992) that could benefit the gesture control. The full metadata format can be found in Appendix A, and some key attributes include:
@@
    Segments: Defines the start and end frames for different gesture phases, including preparation, stroke, and retraction, to support additional features such as gesture holding.
Combining all the data together, when given a gesture command, the module loads the sequence of hand pose data from the gesture database. Downstream modules can use the pose data at each frame as an $21\times 3$ matrix, modify it, and send it to the virtual environment.
@@ 4.3.2. Target Control
The target control module handles object-dependent hand interactions by identifying and aligning the virtual hand with the intended target. It first resolves the target object using positional constraints, such as “below,” “above,” “to the left of,” “to the right of,” “in front of,” “behind,” “closest,” “farthest,” “first,” “last,” “on the left,” “in the middle,” and “on the right.” These constraints are resolved by retrieving the current positions of all virtual objects through a query to Unity application and calculating their relative spatial relationships.
Once the target object has been determined, the module calculates a path to move the hand from its current position to the target object. It first determines the total distance to move, calculates the movement step size by dividing the total distance by the number of frames for this gesture, and cumulatively add step sizes to original hand pose data for all the following gesture frames. This calculation also accounts for the interacting joint (as discussed in Gesture Control) to ensure proper alignment of the hand with the object. For example, if the command is to point at a button, the system uses the interacting joint defined in the gesture metadata to set the index fingertip to align with the center of the button when calculating the movement path. By dynamically modifying the original hand pose sequence, the target control module ensures that gestures interact with objects in the correct fashion.
@@ 4.3.3. Spatial Control
The spatial control module manages the movement and rotation of the virtual hand. It supports translational movements in six directions: “up,” “down,” “left,” “right,” “forward,” and “backward.” Additionally, it enables rotational adjustments, including “pan left,” “pan right,” “roll left,” “roll right,” “tilt up,” and “tilt down.” Specifically, translational movements are applied by adding a speed vector to the hand pose data, while rotational adjustments are calculated by rotating the joint coordinates along the estimated hand axis using the vector of middle metacarpophalangeal joint (MCP) to index MCP, middle MCP to wrist, and the vector perpendicular to these two vectors.
@@ 4.3.4. Temporal Control
Temporal control specifies the timing and state changes of hand interaction, including “stop”, “continue”, “faster”, “slower”, “undo_step”, “redo_step”, and “hold.” Hand pose data from earlier stages is sent to the hand state manager, which keeps a playback speed (frames per second, fps) and playback state (stop, play) that can be adjusted based on temporal control commands. For the hold command, specifically for holding a specific gesture (e.g., as in grab and hold the cube), the temporal control adds a key frame at the interacting frame (as defined in gesture control) and pauses the playback at that frame to hold the gesture.
Additionally, the hand state manager maintains a history of hand poses after each command. This enables “undo” and “redo” functionality, allowing users to revert to previous hand states or redo an action if needed. If an error occurs, the system can quickly restore the last hand state, providing a flexible interaction experience.
At the end, the hand control module outputs the hand pose for the current frame as a list of 63 numbers (21 joints $\times$ (x, y, z)), and sends it to the Unity app to control the virtual hand.
@@ 4.4. Visualization and Feedback
To enhance system transparency and reduce ambiguity, HandProxy uses various visual feedback mechanisms that are overlaid on the user’s viewport. In the prototype Unity app, visualization is implemented as a canvas overlaid on the main camera view. Each visualization is triggered when it receives the command from the HandProxy. Figure 5 shows examples of each feedback type. To ensure users understand how their commands are processed, the system displays both the recognized (full input) and active (which part of the input is being executed) commands. If a command is irrelevant (e.g., not a valid hand interaction) or cannot be interpreted, the system highlights an error message in the recognized input and prompts the user to rephrase the command. For cases where the user refers to an ambiguous object (i.e., multiple similar objects exist in the environment), the system provides disambiguation hints by overlaying numbered labels on each possible target, allowing users to specify the object by its assigned number. To further increase transparency in hand movement execution, the system visualizes the movement path. As shown in Figure 5d, a sequence of arrows appears along the trajectory, indicating the direction in which the hand is moving. If the command involves multiple steps, the full movement sequence is previewed to give users an overview of how the hand will execute the task. These visualizations provide users with greater visibility of the system to bridge the gulf of evaluation, improving both command accuracy and interaction efficiency while reducing errors and misinterpretations.
[…]
4.5. System Latency
The system latency is composed of two main components: speech recognition and processing latency, and command processing latency. Speech recognition and processing latency measures the time from just before an audio chunk is sent via an API request, to when the recognized text is split into commands. Based on 100 requests, the average latency for this process was 0.18 seconds, with a standard deviation of 0.07s. The command processing time is the time it takes for the system to interpret the command and generate a sequence of corresponding hand pose keypoints. We measured such latency for both commands that matched high-priority keywords, and those processed by LLMs. Through a test of 100 commands that matched the high priority keywords, all commands were executed within 0.001 seconds. Through a test of 100 requests that included both simple (e.g., requires one step), complex (e.g., needs to be decomposed into multiple steps), and irrelevant commands (e.g., not a valid interaction instruction), the average response time was 1.39 seconds, with an standard deviation of 0.81s.
4.6. Adapting to New Gestures and Environments
HandProxy can be extended with additional gestures. It accepts hand pose data in the format of a sequence of 21 keypoint coordinates (x, y, z) and a gesture metadata in the format shown in Appendix A. The widely used 21-joint format enables data from various sources to be easily integrated, such as hand pose datasets (Garcia-Hernando et al., 2018), recorded data (e.g., from LeapMotion or MediaPipe), and hand pose generation models (Cha et al., 2024), and HandProxy automatically normalizes the data. Existing built-in controls, such as target, spatial, and temporal controls and other gestures, are still compatible with the newly added gestures.
To adapt to new environments, HandProxy requires the name and optionally descriptive tags of interactable objects, similar to the accessibility metadata for UI elements on mobile (Google, 2024a) or web interfaces (e.g., DOM tree (Mozilla, 2024)). The GPT prompt is dynamically updated with the object information to process new commands. Note that due to limited direct access of the hand input on XR systems, in our current implementation, the virtual hand is added as an object within the application. However, we envision that with system-level support, HandProxy can directly control the virtual hand input of the operating system and retrieve interactable object information to achieve a more unified, cross-application control.
@@ 4.7. Design Iterations
The HandProxy system was iteratively designed and developed through pilot studies with 13 participants recruited from the student group at the authors’ institution. Participants were asked to use HandProxy to perform object selection, manipulation, and transformation tasks in a test virtual environment, and is then interviewed for their experiences and potential improvement suggestions. Here, we summarize key aspects from these iterations, with the goal of providing additional rationale behind our design choices and insights into user preferences.
@@ 4.7.1. Enhancing the Speech Interface
The speech interface is a core component of HandProxy. In our initial design, we followed a traditional turn-based interaction approach, similar to off-the-shelf speech interfaces (app, 2024c), where users issue a command, wait for execution, and then issue the next command. However, our pilot studies revealed that this approach was inefficient for hand control. Many participants preferred to issue multiple commands continuously, describing an entire sequence of actions in one go. The forced pauses between commands disrupted the interaction flow, making it less natural. Additionally, it was difficult for users to issue commands to override current executions, such as stop or undo, as users had to wait for the previous command to finish before issuing a correction. As a result, we redesigned the speech interface to support continuous input, allowing simultaneous speech recognition and action execution. This change also motivated the parallelized system architecture, where each module runs in a separate thread, ensuring that speech processing, command interpretation, and execution run without blocking other processes.
Additionally, we observed variations in user preferences for the level of detail in commands. Some preferred step-by-step instructions, such as “move left,” “grab,” while others issued higher-level commands, such as “put the cube into the basket,” expecting the system to infer intermediate steps. This motivates the choice of integrating LLMs (GPT-4o) than pre-defined commands to support flexible and diverse speech input and leverage their reasoning capabilities to interpret user intent dynamically.
@@ 4.7.2. Integrating Contextual Information
Throughout our iterations, we observed that participants naturally used descriptive references to identify target objects, including shape, appearance, and relative location. Examples include “green ball” or “the object next to the button.” This demonstrated the need for contextual understanding, which motivated our decision to incorporate object metadata into the system.
@@ 4.7.3. Improving System Transparency
Speech recognition errors are a common challenge, and users wanted to know how their commands were interpreted. This feedback motivated several transparency focused design choices, including showing the recognized results and visualizing the path of the expected hand movement. These features ensure that users can intervene in real-time when error happens using commands like stop or redo, which ultimately improves interaction reliability and usability.
@@ 5. User Study
We conducted a user study with 20 participants to evaluate the HandProxy system and gather insights on usage patterns and feedback. Specifically, our research questions are as follows:
@@ 5.1. Participants
We recruited 20 participants (9 female, 11 male, age 18-31) from the student population at our institution. Based on a self-reported survey, 3 participants had no prior experience in immersive environments (e.g., VR/AR), 13 were beginners, and 4 were intermediate users. For experience with hand-gesture-controlled interfaces (e.g., XR headset like Meta Quest or hand tracking cameras like Kinect), 5 participants reported no experience, 12 were beginners, and 3 were intermediate users. Regarding speech interfaces, 1 participant had no experience, 5 were beginners, 12 were intermediate users, and 2 were experts. Examples of speech systems included virtual assistants like Siri or Alexa, ChatGPT voice mode, and speech input on TV remotes. The study was approved by the IRB, and participants received \$30 gift cards as compensation.
@@ 5.2. Apparatus and Procedure
The study sessions were held in a small meeting room. Participants sat in front of a TV connected to a workstation laptop with i9-12900H CPU, which runs the 3D virtual environment in Unity. This setup is to provide an equivalent experience for all participants. As shown in Figure 6, the virtual environment in Unity is designed to have different 3D objects and UI widgets, supporting a variety of interaction tasks. The interactions are triggered only by collision or gesture detection of the virtual hand. The Unity app does not provide API to trigger interaction. Each study session took 2 hours.
@@
    Hand interactions: participants were asked to perform a wide range of interaction tasks for virtual environment (Monteiro et al., 2021), including object selection, manipulation, transformation. The tasks included object dependent gestures (grab {apple, peach, blue cube}, press {confirm, minimize, power} button, pinch {blue cube, volume slider, resize button}, push {confirm, like, power}) and disambiguation (grab the {left, middle, right} watermelon among the 3 watermelons).
@@
    Movement and rotation control: this included hand movement (move {left, up, forward}), object movement (put the {apple, peach, blue cube} into the basket, rotation (turn the brightness knob {clockwise, counterclockwise}). Participants were also asked to perform tasks that required timing control. This included grab the {apple, peach, cube} and move to the left according to a light signal (green = move, red = stop), and to press and hold the {power, like, confirm} button.
@@
5.3. Methodology and Procedure of Data Collection and Analysis